OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x.

OpenStudy (anonymous):

missing something?

OpenStudy (austinl):

Perhaps, the functions?

OpenStudy (debbieg):

And why is this question posted twice? You should delete one of them.

OpenStudy (anonymous):

You should have been given f(x) and g(x)

OpenStudy (anonymous):

I just joined yesterday i'm sorry how would i delete one? and i have a picture

OpenStudy (austinl):

Whoa... that's nasty.

OpenStudy (austinl):

For f(g(x)) you put g(x) in wherever an x is in f(x).

OpenStudy (anonymous):

i suck at inverse problems i get confused easily, mind going step by step?

OpenStudy (austinl):

\[f(g(x)) = \frac{(\frac{-5x-9}{x-1})-9}{(\frac{-5x-9}{x-1})+5}\]

OpenStudy (anonymous):

ok what's the next step

OpenStudy (austinl):

@satellite73 Care to assist?

OpenStudy (debbieg):

Are you sure these are inverses? Or are you supposed to DETERMINE whether they are inverses?

OpenStudy (anonymous):

im supposed to Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x.

OpenStudy (austinl):

Theoretically, you should be able to set f(g(x)) = g(f(x))

OpenStudy (debbieg):

Egads, I thought it wasn't working out when I tried it on paper, but it does. \[f(g(x)) = \dfrac{(\dfrac{-5x-9}{x-1})-9}{(\dfrac{-5x-9}{x-1})+5}\] \[ = \left( \dfrac{(\dfrac{-5x-9}{x-1})-9}{(\dfrac{-5x-9}{x-1})+5} \right)\frac{ x-1 }{x-1}\] \[ = \dfrac{-5x-9-9(x-1)}{-5x-9+5(x-1)} \] \[ = \dfrac{-14x}{-14}=x \]

OpenStudy (debbieg):

Well, that's one direction anyway.

OpenStudy (anonymous):

looks like you are doing fine

OpenStudy (anonymous):

big ugly complex fraction then an orgy of cancellation and you are done will work the same way going the other direction

OpenStudy (austinl):

@satellite73 , couldn't you theoretically set them equal and then cross multiply. Then "simplify".

OpenStudy (anonymous):

no

OpenStudy (austinl):

f(g(x)) = x g(f(x)) = x f(g(x)) = g(f(x)) That doesn't work? And if you get x = x then they are inverse?

OpenStudy (anonymous):

ok i need to process this information, sorry i want to make sure i understand before i just runaway with the end result.this is the easiest way to do it?

OpenStudy (debbieg):

To prove that they are inverses, you have to demonstrate that f(g(x))=x and that g(f(x))=x. That's the only way to do it. :)